Brief Exercise 6-8 (Part Level Submission) At December 31, 2017, the following information was available for E. Hetzel Company: ending inventory $44,000 beginning inventory $58,000, cost of goods sold $260,000, and sales revenue $370,000. Calculate inventory turnover for E. Hetzel Company. (Round answer to 1 decimal place, e.g. 1.5.) Inventory turnover times LINK TO TEXT Attempts: 0 of 3 used SAVE FOR LATER SUBMIT ANSWER ? (b) Calculate days in inventory for E. Hetzel Company. (Round answer to 1 decimal place, e.g. 1.5. Use 365 days for calculation.) Days in inventory s days LINK TO TEXT Attempts: 0 of 3 used SAVE FOR LATER I GUBMTAis
Answer #1

Inventory turnover is an efficiency ratio which calculates the number of times per period a business sells and replaces its entire batch of inventories. It is the ratio of cost of goods sold by a business during an accounting period to the average inventories of the business during the period.

Inventory turnover ratio = Cogs/ average inventory

Cogs = 260000

Average Inventory = (opening Inventory + Ending Inventory)/2

                                       = (58000 + 44000)/ 2

                                       = 102000/ 2

                                       = 51000

Now let's put the values in the formula to calculate Inventory turnover ratio

Inventory turnover ratio = 260000/ 51000

                                                  = 5.1

--------------------------------------------------------------------------------------------------------------------------

Days in inventory = Ending inventory/COGS

                                     = (44000/260000)* 365

                                      = 0.169231* 365

                                         = 61.77 Days

--------------------------------------------------------------------------------------------------------------------------

Feel free to comment if you need further assistance J

Pls rate this answer if you found it useful.